OpenStudy (liizzyliizz):

ok so for the general solution of (x^2-3) dy/dx = xy I got y= (|x^2-3|)/2 +C (is this right or wrong? )

OpenStudy (liizzyliizz):

@amistre64 seem ok? or no?

OpenStudy (amistre64):

\[ (x^2-3) dy/dx = xy\] \[ (x^2-3) dy = xy\ dx\] \[ y^{-1} dy = x (x^2-3)^{-1}\ dx\] \[ ln(y) = \frac12ln(x^2-3)+C\]

OpenStudy (amistre64):

now we get to play with an e :) \[\Large e^{ln(y) = \frac12ln(x^2-3)+C}\] \[y = \large e^{ln(x^2-3)^{1/2}+C}\] \[y = \large (x^2-3)^{1/2}*e^C\] \[y = C(x^2-3)^{1/2}\]

OpenStudy (amistre64):

i think yours got off in the e-ing the sides

OpenStudy (liizzyliizz):

I messed up the e^c at the end.. I was heading in the right direction though lol

OpenStudy (amistre64):

yeah, you were :) practice makes almost perfect tho lol

OpenStudy (liizzyliizz):

the only thing I do not have that you do is the ^1/2 at the end, which has me a little thrown off. o_O

OpenStudy (amistre64):

\[n\ ln(a)=ln(a^n)\] \[e^{ln(a^n)}=a^n\]

OpenStudy (amistre64):

you have to wrap up that 1/2 that pops into the inegration, into the LN to get rid of the LN with an E

OpenStudy (amistre64):

if that makes sense lol

OpenStudy (liizzyliizz):

It somewhat makes sense to me. lol I see what you're doing but i wonder if what I have would be equivalent to your answer lol

OpenStudy (amistre64):

no, since you e-ed out the ln but retained the 1/2 yours is a different beast

OpenStudy (liizzyliizz):

Oh ok I see